### v1

# MINIMIZATION OF NONSMOOTH WEAKLY CONVEX FUNCTION OVER PROX-REGULAR SET FOR ROBUST LOW-RANK MATRIX RECOVERY

Abstract
We propose a prox-regular-type low-rank constrained nonconvex nonsmooth optimization model for Robust Low-Rank Matrix Recovery (RLRMR), i.e., estimate problem of low-rank matrix from an observed signal corrupted by outliers. For RLRMR, the $\ell_{1}$-norm has been utilized as a convex loss to detect outliers as well as to keep tractability of optimization models. Nevertheless, the $\ell_{1}$-norm is not necessarily an ideal robust loss because the $\ell_{1}$-norm tends to overpenalize entries corrupted by outliers of large magnitude. In contrast, the proposed model can employ a weakly convex function as a more robust loss, against outliers, than the $\ell_{1}$-norm. For the proposed model, we present (i) a projected variable smoothing-type algorithm applicable for the minimization of a nonsmooth weakly convex function over a prox-regular set, and (ii) a convergence analysis of the proposed algorithm in terms of stationary point. Numerical experiments demonstrate the effectiveness of the proposed model compared with the existing models that employ the $\ell_{1}$-norm.
Index Terms—  robust low-rank matrix recovery, weak convexity, prox-regularity, nonconvex optimization, variable smoothing
1 Introduction
We consider the task of estimating a low-rank matrix $\bm{X}^{\star}\in\mathbb{R}^{n_{1}\times n_{2}}$ from an observation $\bm{y}\coloneqq\mathcal{A}(\bm{X}^{\star})+\bm{\epsilon}\in\mathbb{R}^{m}$ with a linear operator $\mathcal{A}:\mathbb{R}^{n_{1}\times n_{2}}\to\mathbb{R}^{m}$ and a Gaussian noise $\bm{\epsilon}\in\mathbb{R}^{m}$ [1, 2, 3]. This task arises in, e.g., phase retrieval [4, 5, 6], robust principal component analysis [7, 8], and low-rank matrix completion [9, 10, 11, 12]. In such applications, the observation $\bm{y}$ may be corrupted due to outliers caused by impulse/sparse noise unavoidably in the observation process. In order to mitigate the influence of outliers in the estimation process, Robust Low-Rank Matrix Recovery (RLRMR) has attracted a great attention in the fields of signal processing and machine learning [7, 8, 5, 10, 1, 2, 11, 6, 3]. RLRMR is formulated as follows.
Problem 1.1 (RLRMR).
Let an observation $\bm{y}\in\mathbb{R}^{m}$ satisfy
$(i=1,2,3,\ldots,m)\quad[\bm{y}]_{i}\coloneqq\begin{cases}\mathcal{A}_{i}(\bm{X}^{\star})+[\bm{\epsilon}]_{i},&\mathrm{if}\ i\in\mathcal{I}_{\rm in};\\ \xi_{i},&\mathrm{if}\ i\in\mathcal{I}_{\rm out},\end{cases}\hskip-10.00002pt$ (1)
where $\bm{\epsilon}\in\mathbb{R}^{m}$ is a Gaussian noise, $\mathcal{I}_{\rm in},\mathcal{I}_{\rm out}\subset\{1,2,\ldots,m\}$ denote unknown disjoint index sets of inliers and outliers such that $\mathcal{I}_{\rm in}\cup\mathcal{I}_{\rm out}=\{1,2,\ldots,m\}$ and $\mathcal{I}_{\rm in}\cap\mathcal{I}_{\rm out}=\emptyset$, each $\xi_{i}\in\mathbb{R}\ (i\in\mathcal{I}_{\rm out})$ denotes outlier, and each $\mathcal{A}_{i}:\mathbb{R}^{n_{1}\times n_{2}}\to\mathbb{R}\ (i=1,2,\ldots,m)$ is a known linear operator. For convenience, let $\mathcal{A}:\mathbb{R}^{n_{1}\times n_{2}}\to\mathbb{R}^{m}:\bm{X}\mapsto[\mathcal{A}_{1}(\bm{X})\ \mathcal{A}_{2}(\bm{X})\cdots\mathcal{A}_{m}(\bm{X})]^{\mathsf{T}}$. Then,
[…]
where $\mathcal{L}_{r}\coloneqq\{\bm{X}\in\mathbb{R}^{n_{1}\times n_{2}}\mid\mathrm{Rank}(\bm{X})\leq r\}$ with $r\ll\min\{n_{1},n_{2}\}$.
For Problem 1.1, the convex $\ell_{1}$-norm has been utilized as standard loss to detect outliers [10, 1, 2, 3]. For example, an optimization model with an expression of $\mathcal{L}_{r}=\{\bm{U}\bm{V}^{\mathsf{T}}\in\mathbb{R}^{n_{1}\times n_{2}}\mid\bm{U}\in\mathbb{R}^{n_{1}\times r},\bm{V}\in\mathbb{R}^{n_{2}\times r}\}$ and $\lambda\geq 0$:
$\mathop{\mathrm{minimize}}\limits_{\bm{U}\in\mathbb{R}^{n_{1}\times r},\bm{V}\in\mathbb{R}^{n_{2}\times r}}\left\lVert\bm{y}-\mathcal{A}(\bm{U}\bm{V}^{\mathsf{T}})\right\rVert_{1}+\lambda\left\lVert\bm{U}^{\mathsf{T}}\bm{U}-\bm{V}^{\mathsf{T}}\bm{V}\right\rVert_{F}$ (3)
has been proposed [1, 2] for Problem 1.1, where $\left\lVert\cdot\right\rVert_{F}$ denotes the Frobenius norm, and the second term reduces the scaling ambiguities of $\bm{U}\in\mathbb{R}^{n_{1}\times r}$ and $\bm{V}\in\mathbb{R}^{n_{2}\times r}$ as $\bm{U}\bm{V}^{\mathsf{T}}=(\bm{U}\bm{Q})(\bm{V}\bm{Q}^{-\mathsf{T}})^{\mathsf{T}}$ with invertible matrices $\bm{Q}\in\mathbb{R}^{r\times r}$ (see also (16) for another optimization model). The cost function in (3) enjoys key properties, namely, sharpness and weak convexity [1, Prop. 5, and 6], for applying a subgradient method [13] to the model (3). Theoretical results toward an exact recovery, and numerical experiments in [1] suggest the effectiveness of the model (3) for Problem 1.1.
[…]
Weakly convex functions, e.g., the smoothly clipped absolute deviation (SCAD) [14] (see also (15)) and the Minimax Concave Penalty (MCP) [15], are utilized as promising robust loss functions, where a function $g:\mathbb{R}^{m}\to\mathbb{R}$ is said to be $\eta$-weakly convex if $g(\cdot)+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ is convex with $\eta>0$. Indeed, it has been reported that the use of weakly convex loss improves estimation performance in the literature of robust signal recovery [16, 17]. Hence, the following modification of (3) with a weakly convex loss $\ell:\mathbb{R}\to\mathbb{R}$:
$\hskip-15.00002pt\mathop{\mathrm{minimize}}\limits_{\bm{U}\in\mathbb{R}^{n_{1}\times r},\bm{V}\in\mathbb{R}^{n_{2}\times r}}\sum_{i=1}^{m}\ell\left([\bm{y}]_{i}-\mathcal{A}_{i}(\bm{U}\bm{V}^{\mathsf{T}})\right)+\lambda\left\lVert\bm{U}^{\mathsf{T}}\bm{U}-\bm{V}^{\mathsf{T}}\bm{V}\right\rVert_{F}\hskip-3.00003pt$ (4)
is expected to improve the estimation performance for Problem 1.1. However, for the model (4), any optimization algorithm has not been reported so far mainly because the cost function in (4) with a weakly convex $\ell$ is not necessarily weakly convex.
In this paper, for Problem 1.1, we propose an optimization model with a weakly convex loss $\ell:\mathbb{R}\to\mathbb{R}$:
$\mathop{\mathrm{minimize}}\ \sum_{i=1}^{m}\ell\left([\bm{y}]_{i}-\mathcal{A}_{i}(\bm{X})\right)\ \mathrm{subject\ to\ }\bm{X}\in\mathfrak{L}_{r,\sigma},$ (5)
where $\mathfrak{L}_{r,\sigma}\subset\mathfrak{L}_{r}\subset\mathbb{R}^{n_{1}\times n_{2}}$ with $\sigma>0$ is a closed nonconvex set:
$\hskip-11.99998pt\mathfrak{L}_{r,\sigma}\coloneqq\left\{\bm{X}\in\mathbb{R}^{n_{1}\times n_{2}}\mid\begin{subarray}{c}0<\mathrm{Rank}(\bm{X})\leq r,\\ \sigma\leq\sigma_{j}(\bm{X})\ \mathrm{or}\ \sigma_{j}(\bm{X})=0\ (j=1,2,\ldots,r)\end{subarray}\right\},\hskip-3.99994pt$ (6)
$\sigma_{j}(\cdot)$ stands for the $j$th largest singular value of a given matrix, and $\sigma>0$ works as a lower threshold for nonzero singular values. The set $\mathfrak{L}_{r,\sigma}$ with a small $\sigma>0$ is a reasonable approximation of $\mathfrak{L}_{r}$ because (i) $\mathfrak{L}_{r,\sigma}\cup\{\bm{0}\}$ converges to $\mathfrak{L}_{r}$ as $\sigma\searrow 0$ in the sense of Painlevé-Kuratowski [18, Sect. 4.B]; and (ii) $\mathfrak{L}_{r,\sigma}$ with $\sigma>0$ is prox-regular [19, Thm. 5] (see Problem 1.2 (iii)), which serves as key properties (see, e.g., [20, 21, 18]) regarding the metric projection onto $\mathfrak{L}_{r,\sigma}$, where $\mathfrak{L}_{r}$ is not prox-regular [22]. Moreover, the cost function in (5) with a weakly convex $\ell$ remains weakly convex as a composition of $\ell$ with an affine operator. Hence, the model (5) seems to be more tractable than the model (4). To the best of the authors’ knowledge, this is the first work to employ $\mathcal{L}_{r,\sigma}$ as the constraint set in optimization models for Problem 1.1.
We also present an iterative algorithm for the model (5) via the following optimization problem over a prox-regular set. Indeed, the model (5) is reformulated into Problem 1.2 (see Remark 1.3).
Problem 1.2.
Let $\mathcal{X}$ and $\mathcal{Z}$ be Euclidean spaces. Then,
$\mathrm{find}\ \bm{x}^{\star}\in\mathop{\mathrm{argmin}}\limits_{\bm{x}\in C}F(\bm{x})(\neq\emptyset),$ (7)
where $F\coloneqq g\circ\mathfrak{S}$, $\mathfrak{S}:\mathcal{X}\to\mathcal{Z}$, $g:\mathcal{Z}\to\mathbb{R}$ and $C\subset\mathcal{X}$ satisfy
  1. (i)
    $\mathfrak{S}:\mathcal{X}\to\mathcal{Z}$ is a differentiable mapping such that $\mathfrak{S}$ and its Gâteaux derivative $\mathrm{D}\mathfrak{S}$ are Lipschitz continuous;
  2. (ii)
    $g:\mathcal{Z}\to\mathbb{R}$ is (a) Lipschitz continuous (possibly nonsmooth), (b) $\eta$-weakly convex with $\eta>0$, i.e., $g+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ is convex, and (c) prox-friendly, i.e., $\mathrm{prox}_{\mu g}\ (\mu\in(0,\eta^{-1}))$ (see (9)) is available as a computable tool;
  3. (iii)
    $C\subset\mathcal{X}$ is a nonempty closed prox-regular set [^1] Other than $\mathfrak{L}_{r,\sigma}$ in (6), prox-regular sets include, e.g., closed convex sets, $C^{2}$ embedded submanifolds in $\mathcal{X}$ (see, e.g., [23], and [24, Lemma 2.1]), and $\{\bm{x}\in\mathbb{R}^{n}\mid\left\lVert\bm{x}\right\rVert=1,[\bm{x}]_{i}\geq 0\ \forall i\}$ (see [25, Lemma 3 (ii)]).  [21, Thm. 1.3], i.e., the metric projection $P_{C}:\mathcal{X}\rightrightarrows C:\widebar{\bm{x}}\mapsto\mathop{\mathrm{argmin}}_{\bm{x}\in C}\left\lVert\widebar{\bm{x}}-\bm{x}\right\rVert$ onto $C$ is single-valued on some open subset containing $C$. Moreover, we assume that at least one point $\bm{x}\in P_{C}(\widebar{\bm{x}})$ can be computed for every $\widebar{\bm{x}}\in\mathcal{X}$.
Remark 1.3 (Reformulation of the model (5) into Problem 1.2).
The proposed model (5) is a special instance of Problem 1.2 by setting $\mathcal{X}\coloneqq\mathbb{R}^{n_{1}\times n_{2}}$, $\mathcal{Z}\coloneqq\mathbb{R}^{m}$, $\mathfrak{S}(\bm{X})\coloneqq\bm{y}-\mathcal{A}(\bm{X})$, $g:\mathbb{R}^{m}\to\mathbb{R}:\bm{z}\mapsto\sum_{i=1}^{m}\ell([\bm{z}]_{i})$, and $C\coloneqq\mathfrak{L}_{r,\sigma}$ in (6) with $\sigma>0$, where $\ell$ is Lipschitz continuous, weakly convex and prox-friendly (e.g., the $\ell_{1}$-norm, SCAD [14], and MCP [15]. See also [26, Sect. 2.2] for other examples), and $P_{\mathcal{L}_{r,\sigma}}$ can be computed by using a singular value decomposition (SVD) of a given matrix [19, Cor. 2].
In order to find a stationary point of Problem 1.2, we extend variable smoothing-type algorithms [27, 28, 29, 30] so that even a nonconvex prox-regular set, e.g., $\mathfrak{L}_{r,\sigma}$, can be used as $C$, whereas only a closed convex set $C$ can be used in [27, 28, 29, 30]. The proposed algorithm, named projected variable smoothing, is designed as a projected gradient method with a smoothed surrogate function $^{\mu}g\circ\mathfrak{S}$ of $g\circ\mathfrak{S}$, where $^{\mu}g:\mathcal{Z}\to\mathbb{R}$ with $\mu\in(0,\eta^{-1})$ is the Moreau envelope of $g$ (see (10)). By exploiting its notable properties, e.g., $\lim_{\mu\to 0}{}^{\mu}g(\bm{z})=g(\bm{z})\ (\bm{z}\in\mathcal{Z})$, the proposed algorithm updates $(\bm{x}_{n})_{n=1}^{\infty}\subset C$ as $\bm{x}_{n+1}\in P_{C}(\bm{x}_{n}-\gamma_{n}\nabla({}^{\mu_{n}}g\circ\mathfrak{S})(\bm{x}_{n}))$, where $\gamma_{n}>0$ and $\mu_{n}(\in(0,\eta^{-1}))\searrow 0$ are chosen strategically. We also present an asymptotic convergence analysis, in Theorem 3.4, of the proposed algorithm in terms of a stationary point (see just after (8)).
Numerical experiments demonstrate the effectiveness of proposed model (5) for Problem 1.1 solved by the proposed algorithm.
Related work on Problem 1.2: The theory of prox-regular sets dates back to 1959 [23], and has been studied in, e.g., [20, 21]. However, to the best of our knowledge, only a few recent papers [31, 32] propose iterative algorithms for Problem 1.2 under the assumption that $C$ is proximally smooth [^2] CC is said to be proximally smooth if $P_{C}$ is single-valued over $C+B(\bm{0};\delta)\coloneqq\{\bm{x}+\bm{u}\in\mathcal{X}\mid\bm{x}\in C,\bm{u}\in\mathcal{X},\left\lVert\bm{u}\right\rVert<\delta\}$ with some $\delta>0$. , which is stronger than prox-regularity. Specifically, [31] extends a classical projected subgradient method to the setting of Problem 1.2, and presents its linear-rate convergence under an additional error-bound condition on $F$. The paper [32] extends classical algorithms, e.g., a projected subgradient method and a proximal point method, to the setting of Problem 1.2, and presents iteration-complexity bounds to reach a predetermined tolerance. However, any asymptotic convergence analysis of algorithms in [32] has not been reported yet. In contrast, the proposed algorithm has an asymptotic convergence guarantee (see Thm. 3.4) for Problem 1.2 without error-bound condition required in [31].
Notation: $\mathbb{N}$, $\mathbb{R}$, and $\mathbb{R}_{++}$, denote respectively the sets of all positive integers, all real numbers, and all positive real numbers. $\bm{v}]_{i}\in\mathbb{R}$ denotes $i$th entry of $\bm{v}\in\mathbb{R}^{n}$, For $\widebar{\bm{x}}\in\mathcal{X}$ and a nonempty closed set $E\subset\mathcal{X}$, $\|\widebar{\bm{x}}\|\coloneqq\sqrt{{\left\langle\widebar{\bm{x}},\widebar{\bm{x}}\right\rangle}}$ denotes the Euclidean norm with the standard inner product $\left\langle\cdot,\cdot\right\rangle$}, and $\mathop{\mathrm{dist}}(\widebar{\bm{x}},E)\coloneqq\min\{\left\lVert\bm{v}-\widebar{\bm{x}}\right\rVert\mid\bm{v}\in E\}$ denotes the distance function. For a differentiable mapping $\mathcal{F}:\mathcal{X}\to\mathcal{Z}$, its Gâteaux derivative at $\widebar{\bm{x}}\in\mathcal{X}$ is the linear operator $\mathrm{D}\mathcal{F}(\widebar{\bm{x}}):\mathcal{X}\to\mathcal{Z}:\bm{v}\mapsto\lim_{\mathbb{R}\setminus\{0\}\ni t\to 0}\frac{\mathcal{F}(\widebar{\bm{x}}+t\bm{v})-\mathcal{F}(\widebar{\bm{x}})}{t}$. A mapping $\mathcal{F}:\mathcal{X}\to\mathcal{Z}$ is said to be Lipschitz continuous with a Lipschitz constant $L_{\mathcal{F}}>0$ if $\left\lVert\mathcal{F}(\bm{x}_{1})-\mathcal{F}(\bm{x}_{2})\right\rVert\leq L_{\mathcal{F}}\left\lVert\bm{x}_{1}-\bm{x}_{2}\right\rVert\ (\bm{x}_{1},\bm{x}_{2}\in\mathcal{X})$. For a differentiable function $J:\mathcal{X}\to\mathbb{R}$, $\nabla J(\widebar{\bm{x}})\in\mathcal{X}$ is the gradient of $J$ at $\widebar{\bm{x}}\in\mathcal{X}$ if $\mathrm{D}J(\widebar{\bm{x}})[\bm{v}]={\left\langle\nabla J(\widebar{\bm{x}}),\bm{v}\right\rangle}\ (\bm{v}\in\mathcal{X})$. A function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ is said to be proper if $\mathrm{dom}(J)\coloneqq\{\bm{x}\in\mathcal{X}\mid J(\bm{x})<\infty\}\neq\emptyset$.
2 Preliminary on Nonsmooth Analysis
We review necessary notions and tools in nonsmooth analysis in [18] (see also a comprehensive review paper [33]).
Definition 2.1 (Subdifferential [18, Def. 8.3]).
A vector $\bm{v}\in\mathcal{X}$ is said to be a Fréchet (or regular) subgradient of a proper function $J:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ at $\widebar{\bm{x}}\in\mathrm{dom}(J)$, denoted by $\bm{v}\in\partial_{F}J(\widebar{\bm{x}})$ [18, Def. 8.3], if $\sup\limits_{\epsilon>0}\left(\inf\limits_{0<\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert<\epsilon}\frac{J(\bm{x})-J(\widebar{\bm{x}})-{\left\langle\bm{v},\bm{x}-\widebar{\bm{x}}\right\rangle}}{\left\lVert\bm{x}-\widebar{\bm{x}}\right\rVert}\right)\geq 0$ holds, where $\partial_{F}J$ is called the Fréchet subdifferential of $J$, and $\partial_{F}J(\widebar{\bm{x}})$ at $\widebar{\bm{x}}\notin\mathrm{dom}(J)$ is understood as $\emptyset$. If $J$ is convex, then $\partial_{F}J$ coincides with the convex subdifferential of $J$ [18, Prop. 8.12].
Fermat’s rule [18, Thm. 10.1] serves as a necessary condition:
$\partial_{F}(F+\iota_{C})(\bm{x}^{\star})\ni\bm{0},$ (8)
for local optimality of Problem 1.2, where the indicator function $\iota_{C}:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ is defined as $\iota_{C}(\widebar{\bm{x}})\coloneqq 0$ if $\widebar{\bm{x}}\in C$; $\iota_{C}(\widebar{\bm{x}})\coloneqq+\infty$ if $\widebar{\bm{x}}\notin C$. A point $\bm{x}^{\star}$ enjoying (8) is called a stationary point of Problem 1.2, and finding a stationary point is a reasonable goal for nonconvex optimization [18, 13, 33, 1, 34, 27, 35, 26, 28, 29, 30]. In this paper, we aim to find a stationary point of Problem 1.2.
The proximity operator and the Moreau envelope have been used as computational tools for nonsmooth optimization [36, 37, 34, 27, 35, 26, 28, 29, 30, 6]. For a Euclidean space $\mathcal{H}$, the proximity operator of a proper function $J:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$ with index $\mu>0$ is defined by
$\hskip-5.0pt\mathrm{prox}_{\mu J}:\mathcal{H}\rightrightarrows\mathcal{H}:\widebar{\bm{u}}\mapsto\mathop{\mathrm{argmin}}\limits_{\bm{u}\in\mathcal{H}}\left(J(\bm{u})+\frac{1}{2\mu}\|\bm{u}-\widebar{\bm{u}}\|^{2}\right).$ (9)
By letting $\mathcal{H}\coloneqq\mathcal{X}$ and $J\coloneqq\iota_{C}$ in Problem 1.2, we have the expression $P_{C}=\mathrm{prox}_{\mu\iota_{C}}\ (\mu>0)$, and $P_{C}(\widebar{\bm{x}})\subset C$ is a nonempty and compact set for every $\widebar{\bm{x}}\in\mathcal{X}$ [18, Thm. 1.25]. On the other hand, by letting $\mathcal{H}\coloneqq\mathcal{Z}$ and $J\coloneqq g$ in Problem 1.2, $\mathrm{prox}_{\mu g}(\widebar{\bm{z}})\in\mathcal{Z}$ with $\mu\in(0,\eta^{-1})$ is single-valued for every $\widebar{\bm{z}}\in\mathcal{Z}$ due to the strong convexity of $g+(2\mu)^{-1}\|\cdot-\widebar{\bm{z}}\|^{2}$. By using $\mathrm{prox}_{\mu g}$, the Moreau envelope $^{\mu}g:\mathcal{Z}\to\mathbb{R}\ (\mu\in(0,\eta^{-1}))$ of $g$ is defined by
$^{\mu}g:\mathcal{Z}\to\mathbb{R}:\widebar{\bm{z}}\mapsto g(\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))+\frac{1}{2\mu}\left\lVert\mathrm{prox}_{\mu g}(\widebar{\bm{z}})-\widebar{\bm{z}}\right\rVert^{2}.$ (10)
$^{\mu}g$ has notable properties as an approximation of the nonsmooth function $g$: (i) $\lim_{\mu\to 0}{}^{\mu}g(\widebar{\bm{z}})=g(\widebar{\bm{z}})\ (\widebar{\bm{z}}\in\mathcal{Z})$; (ii) continuous differentiability of $^{\mu}g$ with $\nabla{}^{\mu}g(\widebar{\bm{z}})=\mu^{-1}(\widebar{\bm{z}}-\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))$; (iii) Lipschitz continuity of $\nabla{}^{\mu}g$ (see, e.g, [38, Cor. 3.4]). Both $^{\mu}g$ and $\nabla{}^{\mu}g$ are computable if $g$ is prox-friendly (see Rem. 1.3 for such $g$). Moreover, $\nabla({}^{\mu}g\circ\mathfrak{S})$ enjoys the Lipschitz continuity.
Fact 2.2 (Lipschitzian of $\nabla({}^{\mu}g\circ\mathfrak{S})$ [26, Prop. 4.5(a)]).
Consider $g\circ\mathfrak{S}$ in Problem 1.2. Then, $\nabla({}^{\mu}g\circ\mathfrak{S})$ is Lipschitz continuous with a Lipschitz constant $L_{\nabla({}^{\mu}g\circ\mathfrak{S})}>0$, where there exist $\varpi_{1},\varpi_{2}\in\mathbb{R}_{++}$ such that $L_{\nabla({}^{\mu}g\circ\mathfrak{S})}=\varpi_{1}+\varpi_{2}\mu^{-1}\ (\forall\mu\in(0,(2\eta)^{-1}])$.
3 Projected variable smoothing algorithm
This section presents a key idea, based on a stationarity measure, for finding a stationary point of Problem 1.2 (see (8)), and then an iterative algorithm for Problem 1.2 with a convergence analysis. Due to space limitation, all proofs of the theoretical results in this paper are deferred to an extended manuscript in preparation.
Inspired by [27, 29, 30], we employ a gradient mapping-type stationarity measure $\mathcal{M}_{\gamma}^{F,\iota_{C}}:\mathcal{X}\to\mathbb{R}\ (\gamma\in\mathbb{R}_{++})$ defined as
$(\widebar{\bm{x}}\in\mathcal{X})\quad\mathcal{M}_{\gamma}^{F,\iota_{C}}(\widebar{\bm{x}})\coloneqq\mathop{\mathrm{dist}}\left(\bm{0},\frac{\widebar{\bm{x}}-P_{C}\left(\widebar{\bm{x}}-\gamma\partial_{F}F\left(\widebar{\bm{x}}\right)\right)}{\gamma}\right),$ (11)
with $\frac{\widebar{\bm{x}}-P_{C}\left(\widebar{\bm{x}}-\gamma\partial_{F}F\left(\widebar{\bm{x}}\right)\right)}{\gamma}=\left\{\frac{\widebar{\bm{x}}-\bm{p}}{\gamma}\mid\bm{p}\in P_{C}(\widebar{\bm{x}}-\gamma\bm{v}),\ \bm{v}\in\partial_{F}F(\widebar{\bm{x}})\right\}$. In a special case where $C$ is closed convex, some useful properties of $\mathcal{M}_{\gamma}^{F,\iota_{C}}$ are found in [27, 29, 30]. For example, under the convexity of $C$, a stationary point $\bm{x}^{\star}\in\mathcal{X}$ can be characterized by $\mathcal{M}_{\gamma}^{F,\iota_{C}}(\bm{x}^{\star})=0$ for every $\gamma\in\mathbb{R}_{++}$ [30, Fact III.1 (a)]. Even in the absence of the convexity of $C$, we can characterize a stationary point via $\mathcal{M}_{\gamma}^{F,\iota_{C}}$ for some $\gamma\in\mathbb{R}_{++}$ by Lemma 3.1.
Lemma 3.1 (Stationarity characterization via $\mathcal{M}_{\gamma}^{F,\iota_{C}}$).
$\bm{x}^{\star}\in\mathcal{X}$ is a stationary point of Problem 1.2, i.e., $\partial_{F}(F+\iota_{C})(\bm{x}^{\star})\ni\bm{0}$, if and only if $\mathcal{M}_{\gamma}^{F,\iota_{C}}(\bm{x}^{\star})=0$ holds for some $\gamma\in\mathbb{R}_{++}$.
By Lemma 3.1, we can find a stationary point of Problem 1.2 by finding a point $\bm{x}^{\star}\in\mathcal{X}$ where $\mathcal{M}_{\gamma}^{F,\iota_{C}}(\bm{x}^{\star})=0$ holds for some $\gamma\in\mathbb{R}_{++}$. However, it is still challenging to approximate iteratively a point $\bm{x}^{\star}$ with $\mathcal{M}_{\gamma}^{F,\iota_{C}}(\bm{x}^{\star})=0$ due to the nonsmoothness of $g$ in $F=g\circ\mathfrak{S}$. In contrast, if $F$ were continuously differentiable, then we could exploit powerful arts, e.g., sufficient decrease property, developed for proximal (or projected) gradient method.
Fact 3.2 (Sufficient decrease property, e.g., [39, Lemma 2]).
Consider $C\subset\mathcal{X}$ in Problem 1.2. Let $J:\mathcal{X}\to\mathbb{R}$ be continuously differentiable such that $\nabla J:\mathcal{X}\to\mathcal{X}$ is Lipschitz continuous with a Lipschitz constant $L_{\nabla J}>0$. Let $c\in(0,1/2)$. Then, we have
$(\forall\gamma\in(0,(1-2c)L_{\nabla J}^{-1}],\ \forall\widebar{\bm{x}}\in C,\ \forall\bm{x}\in P_{C}(\widebar{\bm{x}}-\gamma\nabla J(\widebar{\bm{x}}))\subset C)$ (12)
$\ J(\bm{x})\leq J(\widebar{\bm{x}})-c\gamma\left\lVert\frac{\widebar{\bm{x}}-\bm{x}}{\gamma}\right\rVert^{2}\overset{(\clubsuit)}{\leq}J(\widebar{\bm{x}})-c\gamma\left(\mathcal{M}_{\gamma}^{J,\iota_{C}}(\widebar{\bm{x}})\right)^{2},$ (13)
where we used $\left\lVert(\widebar{\bm{x}}-\bm{x})/\gamma\right\rVert\geq\mathcal{M}_{\gamma}^{J,\iota_{C}}(\widebar{\bm{x}})$ (see (11)) in $(\clubsuit)$.
Fact 3.2 motivates us to replace the nonsmooth function $g$ in $\mathcal{M}_{\gamma}^{F,\iota_{C}}=\mathcal{M}_{\gamma}^{g\circ\mathfrak{S},\iota_{C}}$ with its Moreau envelope $^{\mu}g$ in (10). Indeed, the gradient $\nabla({}^{\mu}g\circ\mathfrak{S})$ is Lipschitz continuous by Fact 2.2, and thus $\mathcal{M}_{\gamma}^{{}^{\mu}g\circ\mathfrak{S},\iota_{C}}$ enjoys the sufficient decrease property in (13) with $J\coloneqq{}^{\mu}g\circ\mathfrak{S}$. However, we have a gap between $\mathcal{M}_{\gamma}^{{}^{\mu}g\circ\mathfrak{S},\iota_{C}}$ and $\mathcal{M}_{\gamma}^{F,\iota_{C}}$ because the condition $\mathcal{M}_{\gamma}^{{}^{\mu}g\circ\mathfrak{S},\iota_{C}}(\widebar{\bm{x}})=0$ for $\widebar{\bm{x}}\in\mathcal{X}$ does not imply $\mathcal{M}_{\gamma}^{F,\iota_{C}}(\widebar{\bm{x}})=0$. Theorem 3.3 bridges the gap by passing through an asymptotic behavior of $\mathcal{M}_{\gamma}^{{}^{\mu}g\circ\mathfrak{S},\iota_{C}}$ as $\mu\searrow 0$.
Theorem 3.3 ( [^3] In a special case where $C$ is closed convex and $\gamma_{n}\coloneqq\widebar{\gamma}>0\ (n\in\mathbb{N})$, a similar inequality to (14) is found in our recent paper [30, Thm. III.2 (b)]. However, an extension of [30, Thm. III.2 (b)] to the setting in Thm. 3.3, especially to the variable sequence $(\gamma_{n})_{n=1}^{\infty}$, is non-trivial. Asymptotic property of and $\mathcal{M}_{\gamma}^{{}^{\mu}g\circ\mathfrak{S},\phi}$).
Consider $F=g\circ\mathfrak{S}$ and $C$ in Problem 1.2. Let $(\bm{x}_{n})_{n=1}^{\infty}\subset C$ and $(\gamma_{n})_{n=1}^{\infty}\subset\mathbb{R}_{++}$ converge respectively to some $\widebar{\bm{x}}\in C$ and $\widebar{\gamma}\geq 0$. Then, for $\mu_{n}(\in(0,(2\eta)^{-1}])\searrow 0$ and $F_{n}\coloneqq{}^{\mu_{n}}g\circ\mathfrak{S}\ (n\in\mathbb{N})$,
$\hskip-16.00008pt\liminf_{n\to\infty}\mathcal{M}_{\gamma_{n}}^{F_{n},\iota_{C}}(\bm{x}_{n})\geq\begin{cases}\mathcal{M}_{\widebar{\gamma}}^{F,\iota_{C}}(\widebar{\bm{x}}),&\hskip-8.00003pt\mathrm{if}\ \widebar{\gamma}>0;\\ \mathop{\mathrm{dist}}\left(\bm{0},\partial_{F}(F+\iota_{C})(\widebar{\bm{x}})\right),&\hskip-8.00003pt\mathrm{if}\ \widebar{\gamma}=0\end{cases}\hskip-10.00002pt$ (14)
holds. Moreover, by combining Lemma 3.1 and (8), $\widebar{\bm{x}}$ is a stationary point of Problem 1.2 if $\liminf_{n\to\infty}\mathcal{M}_{\gamma_{n}}^{F_{n},\iota_{C}}(\bm{x}_{n})=0$.
By Thm. 3.3, our goal for finding a stationary point of Problem 1.2 is reduced to finding a sequence $(\bm{x}_{n},\gamma_{n})_{n=1}^{\infty}\subset C\times\mathbb{R}_{++}$ such that $\liminf_{n\to\infty}\mathcal{M}_{\gamma_{n}}^{F_{n},\iota_{C}}(\bm{x}_{n})=0$ with some $\mu_{n}(\in(0,(2\eta)^{-1}])\searrow 0$ and $F_{n}\coloneqq{}^{\mu_{n}}g\circ\mathfrak{S}$. Based on this observation, we present a projected gradient-type method, called a projected variable smoothing algorithm, illustrated in Algorithm 1.
Algorithm 1 Projected variable smoothing for Problem 1.2
[…]
2:for $n=1,2,3,\ldots$ do
3:  Set $\mu_{n}\leftarrow(2\eta)^{-1}n^{-1/\alpha}$ and $F_{n}\coloneqq{}^{\mu_{n}}g\circ\mathfrak{S}$
4:  for $m=0,1,2,\ldots$ do
5:   $\bm{x}\in P_{C}(\bm{x}_{n}-\rho^{m}\widetilde{\gamma}\nabla F_{n}(\bm{x}_{n}))\subset C$
6:   if $F_{n}(\bm{x})\leq F_{n}(\bm{x}_{n})-c\rho^{m}\widetilde{\gamma}\left\lVert\frac{\bm{x}_{n}-\bm{x}}{\rho^{m}\widetilde{\gamma}}\right\rVert^{2}$ holds then
[…]
8:   end if
9:  end for
10:end for
In Alg. 1, we update estimates as $\bm{x}_{n+1}\in P_{C}(\bm{x}_{n}-\gamma_{n}\nabla F_{n}(\bm{x}_{n}))$ with a stepsize $\gamma_{n}$ such that the sufficient decrease condition in (13) with $\widebar{\bm{x}}\coloneqq\bm{x}_{n}$ and $J\coloneqq{}^{\mu_{n}}g\circ\mathfrak{S}$ is achieved by $(\bm{x},\gamma)\coloneqq(\bm{x}_{n+1},\gamma_{n})$, where $\mu_{n}\coloneqq(2\eta)^{-1}n^{-1/\alpha}$ with $\alpha\geq 1$ (Note: $\alpha=3$ is recommended in the literature of variable smoothing-type algorithms [34, 35, 26, 27, 28, 6, 30] that can not be applied to Problem 1.2 due to the nonconvexity of $C$). To find such a pair $(\bm{x}_{n+1},\gamma_{n})$, we employ a standard backtracking algorithm in line 4-9 of Alg. 1 (see, e.g., [26, 30]). Thanks to Fact 3.2 with Fact 2.2, we can find such $(\bm{x}_{n+1},\gamma_{n})$ in at most $\max\left\{1,\left\lceil\log_{\rho}\left(\frac{(1-2c)}{\widetilde{\gamma}L_{\nabla({}^{\mu_{n}}g\circ\mathfrak{S})}}\right)\right\rceil\right\}$ backtracking steps in line 4-9, where Alg. 1 does not require any knowledge on a constant $L_{\nabla({}^{\mu_{n}}g\circ\mathfrak{S})}=\varpi_{1}+\varpi_{2}\mu_{n}^{-1}$ in Fact 2.2.
We present a convergence analysis of Alg. 1 in Theorem 3.4444To show Theorem 3.4, we can apply the discussion similar to the proof of [30, Thm. III.6 with $\phi\coloneqq\iota_{C}$] although $\phi$ in [30] is assumed to be convex (Note: $\iota_{C}$ (see after (8)) with $C$ in Problem 1.2 is nonconvex). .
Theorem 3.4 (Convergence analysis of Alg. 1).
Consider Problem 1.2. Choose $\bm{x}_{1}\in C$, $c\in(0,2^{-1})$, $\widetilde{\gamma}\in\mathbb{R}_{++}$, and $\alpha\geq 1$. For $(\bm{x}_{n})_{n=1}^{\infty}\subset C$ generated by Alg. 1, the following hold with $F_{n}={}^{\mu_{n}}g\circ\mathfrak{S}\ (n\in\mathbb{N})$ and $\mu_{n}\coloneqq(2\eta)^{-1}n^{-1/\alpha}$:
  1. (a)
    $\liminf_{n\to\infty}\mathcal{M}_{\gamma_{n}}^{F_{n},\iota_{C}}(\bm{x}_{n})=0$}.
  2. (b)
    Choose a subsequence $(\bm{x}_{m(l)})_{l=1}^{\infty}\subset\mathcal{X}$ of $(\bm{x}_{n})_{n=1}^{\infty}$ such that $\lim_{l\to\infty}\mathcal{M}_{\gamma_{m(l)}}^{F_{m(l)},\iota_{C}}(\bm{x}_{m(l)})=0$, where $m:\mathbb{N}\to\mathbb{N}$ is monotonically increasing. Then, every cluster point $\bm{x}^{\star}\in C$ of $(\bm{x}_{m(l)})_{l=1}^{\infty}$ is a stationary point of Problem 1.2.
Remark 3.5 (Extension of Alg. 1 to the minimization of $F+\phi$ with a prox-regular function $\phi$).
For simplicity, we are focusing on Problem 1.2 in this paper. Nevertheless, Lemma 3.1, and Thms. 3.3 and 3.4 can be extended to a case where $\iota_{C}$ is replaced by a proper lower semicontinuous and prox-regular[18, Def. 13.27] function $\phi:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ such that $\partial_{F}\phi$ is outer semicontinuous [18, Def. 5.4] at every $\widebar{\bm{x}}\in\mathrm{dom}(\phi)$ (Note: such a $\phi$ includes, e.g., the $\ell_{0}$-pseudonorm). In this case, the measure in (11) is extended to $\mathcal{M}_{\gamma}^{F,\phi}(\widebar{\bm{x}})\coloneqq\inf\left\{\left\lVert\frac{\widebar{\bm{x}}-\bm{p}}{\gamma}\right\rVert\mid\bm{p}\in\mathrm{prox}_{\gamma\phi}(\widebar{\bm{x}}-\gamma\bm{v}),\ \bm{v}\in\partial_{F}F(\widebar{\bm{x}})\right\}$. To the problem for finding a stationary point of $F+\phi$, we can apply a modified version of Alg. 1, in the sense of Thm. 3.4, by replacing (i) $P_{C}$ in line 5 with $\mathrm{prox}_{\rho^{m}\widetilde{\gamma}}$; and (ii) the condition in line 6 with the condition $(F_{n}+\phi)(\bm{x})\leq(F_{n}+\phi)(\bm{x}_{n})-c\rho^{m}\widetilde{\gamma}\left\lVert\frac{\bm{x}_{n}-\bm{x}}{\rho^{m}\widetilde{\gamma}}\right\rVert$, where $\widetilde{\gamma}\in\mathbb{R}_{++}$ is chosen such that $\phi+\frac{1}{2\widetilde{\gamma}}\left\lVert\cdot\right\rVert^{2}$ is bounded below.
4 Numerical Experiments on Synthetic Data
In a scenario of RLRMR in Problem 1.1, we demonstrate the performance of the proposed model (5) with $\sigma\coloneqq 1$ for $\mathcal{L}_{r,\sigma}$ in (6) and with two loss functions: (i) the $\ell_{1}$-norm $\left\lVert\cdot\right\rVert_{1}$, i.e., $\ell\coloneqq\lvert\cdot\rvert$, and (ii) SCAD [14], i.e., $\ell\coloneqq r^{\rm SCAD}_{\theta}$ with a parameter $\theta>2$:
$r^{\rm SCAD}_{\theta}:\mathbb{R}\to\mathbb{R}:t\mapsto\begin{cases}\lvert t\rvert,&\mathrm{if}\ \lvert t\rvert\leq 1;\\ \frac{-t^{2}+2\theta\lvert t\rvert-1}{2(\theta-1)},&\mathrm{if}\ 1<\lvert t\rvert\leq\theta;\\ \frac{\theta+1}{2},&\mathrm{if}\ \lvert t\rvert>\theta,\end{cases}$ (15)
which is $(\theta-1)^{-1}$-weakly convex (e.g., [34]). For comparisons, we employed two state-of-the-art models: the model (3) in [1] and the model in [3] with weights $\lambda,\beta\in\mathbb{R}_{++}$:
$\mathop{\mathrm{minimize}}_{\bm{X}\in\mathbb{R}^{n_{1}\times n_{2}}}\ \left\lVert\bm{y}-\mathcal{A}(\bm{X})\right\rVert_{1}+\lambda(\left\lVert\bm{X}\right\rVert_{\rm nuc}-\beta\left\lVert\bm{X}\right\rVert_{F}),$ (16)
where $\left\lVert\cdot\right\rVert_{\rm nuc}$ denotes the nuclear norm, and the second term in (16) is a regularizer to promote the low-rankness of $\bm{X}$. We applied (a) Alg. 1 with $(c,\rho,\widetilde{\gamma},\alpha)\coloneqq(2^{-13},2^{-1},1,3)$ to (5) (see also Remark 1.3); (b) a subgradient method [1, Alg. 2.1] (see also [13]) to (3); and (c) a difference-of-convex algorithm (DCA) [3, Alg. 1] (see also [40]) to (16), where an ADMM, e.g., [41], was employed for a convex subproblem in DCA [3, Alg. 1]. All algorithms were terminated when runtime exceeded $60$ (s) or the relative difference $\lvert J(\bm{x}_{n+1})-J(\bm{x}_{n})\rvert/\lvert J(\bm{x}_{n})\rvert$ was less than $10^{-9}$, where $J$ is the cost function in (3), (5), or (16), and $\bm{x}_{n}$ is an $n$th estimate of a solution generated by each algorithm. All experiments were performed by MATLAB on MacBookPro (Apple M3, 16 GB).
For every trial, according to [1, 3], we generated $\bm{X}^{\star}\coloneqq\bm{U}^{\star}\bm{V}^{\star\mathsf{T}}\in\mathcal{L}_{r}\subset\mathbb{R}^{n_{1}\times n_{2}}$ with random $\bm{U}^{\star}\in\mathbb{R}^{n_{1}\times r}$ and $\bm{V}^{\star}\in\mathbb{R}^{n_{2}\times r}$, $\mathcal{A}_{i}:\mathbb{R}^{n_{1}\times n_{2}}\to\mathbb{R}\ (i=1,2,\ldots,m)$, and $\bm{y}\in\mathbb{R}^{m}$ as follows. Each entry of $\bm{U}^{\star}$, $\bm{V}^{\star}$ and all $\mathcal{A}_{i}$, as matrix expressions of $\mathcal{A}_{i}$, was sampled from the normal distribution $\mathcal{N}(0,1)$. We generated $\bm{y}\in\mathbb{R}^{m}$ by (1), where each entry of noise $\bm{\epsilon}\in\mathbb{R}^{m}$ was sampled from $\mathcal{N}(0,10^{-6})$ according to [3], the index set $\mathcal{I}_{\rm out}$ of outliers was randomly chosen, and each outlier $\xi_{i}\in\mathbb{R}\ (i\in\mathcal{I}_{\rm out})$ was generated with $\Omega\coloneqq\max_{i=1,2,\ldots,m}{\lvert\mathcal{A}_{i}(\bm{X}^{\star})\rvert}$ from (i) uniform distribution $\mathcal{U}_{[-\Omega,\Omega]}$; and (ii) Cauchy distribution, i.e., $\xi_{i}\coloneqq\Omega\tan(\pi u_{i}/2)$ with $u_{i}\in(-1,1)$ sampled from $\mathcal{U}_{(-1,1)}$.
Table 1 demonstrates the averaged RMSE and runtime of each model [^5] For (5) with $\ell\coloneqq r_{\theta}^{\rm SCAD}$ in (15), we employed $\theta\in\{2.5,2.7,2.9,3.1\}$. For (3), $\lambda\in\{10^{-2},10^{-1},1,2,5\}$, was chosen. Moreover, stepsize of the subgradient method used in [1] at $n$th iteration was chosen from $0.95^{n}$ or $2\times 0.95^{n}$, where the first choice was suggested by [1] while the second choice achieved better results in some cases in our scenario. For (16), $\lambda\coloneqq t\sqrt{mn_{2}\log(n_{1}+n_{2})}$ with $t\in\{0.1,0,5\}$ and $\beta\in\{0.1,0.5,0.9\}$ were chosen according to [3]. over $100$ trials with $(n_{1},n_{2},r)=(40,50,5)$, $m=p_{m}n_{1}n_{2}$ and $\lvert\mathcal{I}_{\rm out}\rvert=p_{\lvert\mathcal{I}_{\rm out}\rvert}m$ for $p_{m}\in\{0.3,0.5,0.7\}$ and $p_{\lvert\mathcal{I}_{\rm out}\rvert}\in\{0.7,0.5\}$, where we used RMSE $\lVert\widehat{\bm{X}}-\bm{X}^{\star}\rVert_{F}/\sqrt{n_{1}n_{2}}$ to measure a recovery error with the final estimate $\widehat{\bm{X}}\in\mathbb{R}^{n_{1}\times n_{2}}$ of algorithm. We note that Problem 1.1 becomes challenging as $p_{m}$ is smaller and $p_{\lvert\mathcal{I}_{\rm out}\rvert}$ is larger.
From Table 1, we observe that convergence speeds of the proposed models (5) and the model (16) are much slower than that of the model (3). This is because singular value decompositions (SVD) of $n_{1}$-by-$n_{2}$ matrices are required at every iteration in solvers for (5) (see also Remark 1.3) and for (16) while any SVD is not required in a subgradient method for (3) because of the decomposition $\bm{X}=\bm{U}\bm{V}^{\mathsf{T}}$ in (3) (Note: acceleration of the proposed Alg. 1 is beyond the scope of this paper, and will be addressed in future work). Moreover, from Table 1, the proposed model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ achieves smaller RMSE than the model (5) with $\ell\coloneqq\lvert\cdot\rvert$ in a shorter runtime; and RMSE of the existing models (3) and (16) are much higher than that of the model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ in particular for $p_{m}\in\{0.3,0.5\}$. These observations imply that a weakly convex $r^{\rm SCAD}_{\theta}$ is more robust against outliers, at least in the proposed model (5), than a convex $\lvert\cdot\rvert$ as we expected; and the proposed model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ outperforms the existing models (3) and (16) in terms of recovery accuracy.
Table 1: Averaged RMSE and (averaged runtime [s]) for Problem 1.1
Setting Proposed model (5) Existing models
[…]
$(0.3,0.7)$ uniform 6.81E-01 (59.9) 3.46E-05 (21.8) 9.79E-01 (0.7) 1.89E-01 (53.3)
Cauchy 6.56E-01 (59.8) 3.47E-05 (23.6) 9.93E-01 (0.6) 4.59E-01 (58.7)
$(0.3,0.5)$ uniform 7.02E-01 (60.0) 3.83E-05 (20.2) 1.04E+00 (0.7) 1.83E-01 (55.9)
Cauchy 6.87E-01 (60.0) 3.81E-05 (20.3) 1.10E+00 (0.7) 1.69E-01 (55.8)
$(0.5,0.7)$ uniform 5.09E-05 (58.4) 1.92E-05 (5.5) 2.92E-02 (1.1) 4.78E-03 (59.3)
Cauchy 5.47E-05 (50.9) 1.91E-05 (6.9) 7.39E-03 (1.2) 4.53E-03 (58.5)
$(0.5,0.5)$ uniform 5.02E-05 (57.7) 1.92E-05 (6.2) 8.11E-03 (1.0) 5.52E-03 (59.3)
Cauchy 5.48E-05 (50.7) 1.98E-05 (5.7) 3.61E-03 (1.1) 5.05E-03 (58.0)
$(0.7,0.7)$ uniform 3.17E-05 (58.6) 1.50E-05 (4.4) 1.72E-05 (1.3) 1.35E-03 (7.4)
Cauchy 3.34E-05 (52.8) 1.48E-05 (4.0) 1.74E-05 (1.6) 1.88E-03 (9.2)
$(0.7,0.5)$ uniform 3.15E-05 (58.2) 1.48E-05 (4.0) 1.75E-05 (1.7) 1.31E-03 (8.4)
Cauchy 3.18E-05 (55.9) 1.51E-05 (5.1) 1.74E-05 (1.4) 1.97E-03 (9.1)
5 Concluding Remarks
In this paper, we proposed a prox-regular-type low-rank constrained optimization model, incorporating a weakly convex loss function, for robust low-rank matrix recovery. For the proposed model, we presented an optimization algorithm, for minimization of a weakly convex function over a prox-regular set, with guaranteed convergence in terms of stationary point. Our numerical experiments demonstrate that the proposed model with a weakly convex loss function, called SCAD, dramatically improves estimation performance of robust low-rank matrix recovery under severe outliers. Finally, we defer further numerical evaluations and complete proofs, including the extensions discussed in Remark 3.5, to an extended manuscript in preparation.

### v2

@@
 where $\mathcal{L}_{r}\coloneqq\{\bm{X}\in\mathbb{R}^{n_{1}\times n_{2}}\mid\mathrm{Rank}(\bm{X})\leq r\}$ with $r\ll\min\{n_{1},n_{2}\}$.
-For Problem 1.1, the convex $\ell_{1}$-norm has been utilized as standard loss to detect outliers [10, 1, 2, 3].
+For Problem 1.1, the convex $\ell_{1}$-norm $\left\lVert\cdot\right\rVert_{1}$ has been utilized as standard loss to detect outliers [10, 1, 2, 3].
 For example, an optimization model with an expression of $\mathcal{L}_{r}=\{\bm{U}\bm{V}^{\mathsf{T}}\in\mathbb{R}^{n_{1}\times n_{2}}\mid\bm{U}\in\mathbb{R}^{n_{1}\times r},\bm{V}\in\mathbb{R}^{n_{2}\times r}\}$ and $\lambda\geq 0$:
@@
 is expected to improve the estimation performance for Problem 1.1.
-However, for the model (4), any optimization algorithm has not been reported so far mainly because the cost function in (4) with a weakly convex $\ell$ is not necessarily weakly convex.
+However, for the model (4), any reliable optimization algorithm has not been reported so far mainly because the cost function in (4) with a weakly convex $\ell$ is not necessarily weakly convex.
 In this paper, for Problem 1.1, we propose an optimization model with a weakly convex loss $\ell:\mathbb{R}\to\mathbb{R}$:
@@
   2. (ii)
-    $g:\mathcal{Z}\to\mathbb{R}$ is (a) Lipschitz continuous (possibly nonsmooth), (b) $\eta$-weakly convex with $\eta>0$, i.e., $g+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ is convex, and (c) prox-friendly, i.e., $\mathrm{prox}_{\mu g}\ (\mu\in(0,\eta^{-1}))$ (see (9)) is available as a computable tool;
+    $g:\mathcal{Z}\to\mathbb{R}$ is (a) $L_{g}$-Lipschitz continuous with $L_{g}>0$ (possibly nonsmooth), (b) $\eta$-weakly convex with $\eta>0$, i.e., $g+\frac{\eta}{2}\left\lVert\cdot\right\rVert^{2}$ is convex, and (c) prox-friendly, i.e., $\mathrm{prox}_{\mu g}\ (\mu\in(0,\eta^{-1}))$ (see (9)) is available as a computable tool;
   3. (iii)
@@
 [21, Thm.
-1.3], i.e., the metric projection $P_{C}:\mathcal{X}\rightrightarrows C:\widebar{\bm{x}}\mapsto\mathop{\mathrm{argmin}}_{\bm{x}\in C}\left\lVert\widebar{\bm{x}}-\bm{x}\right\rVert$ onto $C$ is single-valued on some open subset containing $C$.
+1.3], i.e., the metric projection $P_{C}:\mathcal{X}\rightrightarrows C:\widebar{\bm{x}}\mapsto\mathop{\mathrm{argmin}}_{\bm{x}\in C}\left\lVert\widebar{\bm{x}}-\bm{x}\right\rVert$ onto $C$ is single-valued on some open superset of $C$.
 Moreover, we assume that at least one point $\bm{x}\in P_{C}(\widebar{\bm{x}})$ can be computed for every $\widebar{\bm{x}}\in\mathcal{X}$.
@@
 In this paper, we aim to find a stationary point of Problem 1.2.
-The proximity operator and the Moreau envelope have been used as computational tools for nonsmooth optimization [36, 37, 34, 27, 35, 26, 28, 29, 30, 6].
+The proximity operator and the Moreau envelope have been used as computational tools for nonsmooth optimization [36, 34, 27, 35, 26, 28, 29, 30, 6].
 For a Euclidean space $\mathcal{H}$, the proximity operator of a proper function $J:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$ with index $\mu>0$ is defined by
@@
 $^{\mu}g:\mathcal{Z}\to\mathbb{R}:\widebar{\bm{z}}\mapsto g(\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))+\frac{1}{2\mu}\left\lVert\mathrm{prox}_{\mu g}(\widebar{\bm{z}})-\widebar{\bm{z}}\right\rVert^{2}.$ (10)
-$^{\mu}g$ has notable properties as an approximation of the nonsmooth function $g$: (i) $\lim_{\mu\to 0}{}^{\mu}g(\widebar{\bm{z}})=g(\widebar{\bm{z}})\ (\widebar{\bm{z}}\in\mathcal{Z})$; (ii) continuous differentiability of $^{\mu}g$ with $\nabla{}^{\mu}g(\widebar{\bm{z}})=\mu^{-1}(\widebar{\bm{z}}-\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))$; (iii) Lipschitz continuity of $\nabla{}^{\mu}g$ (see, e.g, [38, Cor.
+$^{\mu}g$ has notable properties as an approximation of the nonsmooth function $g$: (i) $\lim_{\mu\to 0}{}^{\mu}g(\widebar{\bm{z}})=g(\widebar{\bm{z}})\ (\widebar{\bm{z}}\in\mathcal{Z})$; (ii) continuous differentiability of $^{\mu}g$ with $\nabla{}^{\mu}g(\widebar{\bm{z}})=\mu^{-1}(\widebar{\bm{z}}-\mathrm{prox}_{\mu g}(\widebar{\bm{z}}))$; (iii) Lipschitz continuity of $\nabla{}^{\mu}g$ (see, e.g, [37, Cor.
 3.4]).
@@
 Moreover, $\nabla({}^{\mu}g\circ\mathfrak{S})$ enjoys the Lipschitz continuity.
-Fact 2.2 (Lipschitzian of $\nabla({}^{\mu}g\circ\mathfrak{S})$ [26, Prop. 4.5(a)]).
+Fact 2.2 (Lipschitzian of $\nabla({}^{\mu}g\circ\mathfrak{S})$ [26, Prop. 4.2(a)]).
 Consider $g\circ\mathfrak{S}$ in Problem 1.2.
@@
 In contrast, if $F$ were continuously differentiable, then we could exploit powerful arts, e.g., sufficient decrease property, developed for proximal (or projected) gradient method.
-Fact 3.2 (Sufficient decrease property, e.g., [39, Lemma 2]).
+Fact 3.2 (Sufficient decrease property, e.g., [38, Lemma 2]).
 Consider $C\subset\mathcal{X}$ in Problem 1.2.
@@
 $(\forall\gamma\in(0,(1-2c)L_{\nabla J}^{-1}],\ \forall\widebar{\bm{x}}\in C,\ \forall\bm{x}\in P_{C}(\widebar{\bm{x}}-\gamma\nabla J(\widebar{\bm{x}}))\subset C)$ (12)
-$\ J(\bm{x})\leq J(\widebar{\bm{x}})-c\gamma\left\lVert\frac{\widebar{\bm{x}}-\bm{x}}{\gamma}\right\rVert^{2}\overset{(\clubsuit)}{\leq}J(\widebar{\bm{x}})-c\gamma\left(\mathcal{M}_{\gamma}^{J,\iota_{C}}(\widebar{\bm{x}})\right)^{2},$ (13)
-where we used $\left\lVert(\widebar{\bm{x}}-\bm{x})/\gamma\right\rVert\geq\mathcal{M}_{\gamma}^{J,\iota_{C}}(\widebar{\bm{x}})$ (see (11)) in $(\clubsuit)$.
+$\ J(\bm{x})\leq J(\widebar{\bm{x}})-c\gamma\left\lVert(\widebar{\bm{x}}-\bm{x})/\gamma\right\rVert^{2}\leq J(\widebar{\bm{x}})-c\gamma\left(\mathcal{M}_{\gamma}^{J,\iota_{C}}(\widebar{\bm{x}})\right)^{2}.$ (13)
 Fact 3.2 motivates us to replace the nonsmooth function $g$ in $\mathcal{M}_{\gamma}^{F,\iota_{C}}=\mathcal{M}_{\gamma}^{g\circ\mathfrak{S},\iota_{C}}$ with its Moreau envelope $^{\mu}g$ in (10).
@@
 Then, for $\mu_{n}(\in(0,(2\eta)^{-1}])\searrow 0$ and $F_{n}\coloneqq{}^{\mu_{n}}g\circ\mathfrak{S}\ (n\in\mathbb{N})$,
-$\hskip-16.00008pt\liminf_{n\to\infty}\mathcal{M}_{\gamma_{n}}^{F_{n},\iota_{C}}(\bm{x}_{n})\geq\begin{cases}\mathcal{M}_{\widebar{\gamma}}^{F,\iota_{C}}(\widebar{\bm{x}}),&\hskip-8.00003pt\mathrm{if}\ \widebar{\gamma}>0;\\ \mathop{\mathrm{dist}}\left(\bm{0},\partial_{F}(F+\iota_{C})(\widebar{\bm{x}})\right),&\hskip-8.00003pt\mathrm{if}\ \widebar{\gamma}=0\end{cases}\hskip-10.00002pt$ (14)
+$\hskip-16.00008pt\liminf_{n\to\infty}\mathcal{M}_{\gamma_{n}}^{F_{n},\iota_{C}}(\bm{x}_{n})\geq\begin{cases}\mathcal{M}_{\widebar{\gamma}}^{F,\iota_{C}}(\widebar{\bm{x}}),&\hskip-8.00003pt\mathrm{if}\ \widebar{\gamma}>0;\\ \mathop{\mathrm{dist}}\left(\bm{0},\partial_{F}(F+\iota_{C})(\widebar{\bm{x}})\right),&\hskip-8.00003pt\par\mathrm{if}\ \widebar{\gamma}=0\end{cases}\hskip-10.00002pt$ (14)
 holds. Moreover, by combining Lemma 3.1 and (8), $\widebar{\bm{x}}$ is a stationary point of Problem 1.2 if $\liminf_{n\to\infty}\mathcal{M}_{\gamma_{n}}^{F_{n},\iota_{C}}(\bm{x}_{n})=0$.
@@
 In Alg.
-1, we update estimates as $\bm{x}_{n+1}\in P_{C}(\bm{x}_{n}-\gamma_{n}\nabla F_{n}(\bm{x}_{n}))$ with a stepsize $\gamma_{n}$ such that the sufficient decrease condition in (13) with $\widebar{\bm{x}}\coloneqq\bm{x}_{n}$ and $J\coloneqq{}^{\mu_{n}}g\circ\mathfrak{S}$ is achieved by $(\bm{x},\gamma)\coloneqq(\bm{x}_{n+1},\gamma_{n})$, where $\mu_{n}\coloneqq(2\eta)^{-1}n^{-1/\alpha}$ with $\alpha\geq 1$ (Note: $\alpha=3$ is recommended in the literature of variable smoothing-type algorithms [34, 35, 26, 27, 28, 6, 30] that can not be applied to Problem 1.2 due to the nonconvexity of $C$).
+1, we update estimates as $\bm{x}_{n+1}\in P_{C}(\bm{x}_{n}-\gamma_{n}\nabla F_{n}(\bm{x}_{n}))$ with a stepsize $\gamma_{n}$ such that the sufficient decrease condition in (13) with $\widebar{\bm{x}}\coloneqq\bm{x}_{n}$ and $J\coloneqq{}^{\mu_{n}}g\circ\mathfrak{S}$ is achieved by $(\bm{x},\gamma)\coloneqq(\bm{x}_{n+1},\gamma_{n})$, where $\mu_{n}\coloneqq(2\eta)^{-1}n^{-1/\alpha}$ with $\alpha\geq 1$ (Note: $\alpha=3$ is recommended in the literature of variable smoothing-type algorithms [34, 35, 26, 27, 28, 29, 6, 30] that can not be applied to Problem 1.2 due to the nonconvexity of $C$).
 To find such a pair $(\bm{x}_{n+1},\gamma_{n})$, we employ a standard backtracking algorithm in line 4-9 of Alg.
 1 (see, e.g., [26, 30]).
-Thanks to Fact 3.2 with Fact 2.2, we can find such $(\bm{x}_{n+1},\gamma_{n})$ in at most $\max\left\{1,\left\lceil\log_{\rho}\left(\frac{(1-2c)}{\widetilde{\gamma}L_{\nabla({}^{\mu_{n}}g\circ\mathfrak{S})}}\right)\right\rceil\right\}$ backtracking steps in line 4-9, where Alg.
+Thanks to Fact 3.2 with Fact 2.2, we can find such $(\bm{x}_{n+1},\gamma_{n})$ in at most $\max\left\{1,\left\lceil\log_{\rho}\left(\frac{1-2c}{\widetilde{\gamma}L_{\nabla({}^{\mu_{n}}g\circ\mathfrak{S})}}\right)\right\rceil\right\}$ backtracking steps in line 4-9, where Alg.
 1 does not require any knowledge on a constant $L_{\nabla({}^{\mu_{n}}g\circ\mathfrak{S})}=\varpi_{1}+\varpi_{2}\mu_{n}^{-1}$ in Fact 2.2.
-We present a convergence analysis of Alg.
-1 in Theorem 3.4444To show Theorem 3.4, we can apply the discussion similar to the proof of [30, Thm.
-III.6 with $\phi\coloneqq\iota_{C}$] although $\phi$ in [30] is assumed to be convex (Note: $\iota_{C}$ (see after (8)) with $C$ in Problem 1.2 is nonconvex).
-.
+Finally, we present a convergence analysis of Alg.
+1 in Theorem 3.4.
 Theorem 3.4 (Convergence analysis of Alg. 1).
 Consider Problem 1.2.
-Choose $\bm{x}_{1}\in C$, $c\in(0,2^{-1})$, $\widetilde{\gamma}\in\mathbb{R}_{++}$, and $\alpha\geq 1$.
+Choose $\bm{x}_{1}\in C$, $c\in(0,2^{-1})$, $\rho\in(0,1)$, $\widetilde{\gamma}\in\mathbb{R}_{++}$, and $\alpha\geq 1$.
 For $(\bm{x}_{n})_{n=1}^{\infty}\subset C$ generated by Alg.
@@
   1. (a)
+    If $\alpha>1$, then $\min\limits_{1\leq n\leq k}\mathcal{M}_{\gamma_{n}}^{F_{n},\iota_{C}}(\bm{x}_{n})\leq\sqrt{\frac{\chi}{(k+1)^{1-\alpha^{-1}}-1}}$} $(k\in\mathbb{N})$ holds with $\chi\coloneqq\frac{\left(F(\bm{x}_{1})-\inf_{\bm{x}\in C}F(\bm{x})+L_{g}^{2}(2\eta)^{-1}\right)(\varpi_{1}+2\eta\varpi_{2})}{2\eta c\min\left\{\widetilde{\gamma}(\varpi_{1}+2\eta\varpi_{2}),\rho(1-2c)\right\}}\in\mathbb{R}_{++}$, where $\varpi_{1},\varpi_{2}\in\mathbb{R}_{++}$ are given in Fact 2.2.
+  2. (b)
     $\liminf_{n\to\infty}\mathcal{M}_{\gamma_{n}}^{F_{n},\iota_{C}}(\bm{x}_{n})=0$}.
-  2. (b)
-    Choose a subsequence $(\bm{x}_{m(l)})_{l=1}^{\infty}\subset\mathcal{X}$ of $(\bm{x}_{n})_{n=1}^{\infty}$ such that $\lim_{l\to\infty}\mathcal{M}_{\gamma_{m(l)}}^{F_{m(l)},\iota_{C}}(\bm{x}_{m(l)})=0$, where $m:\mathbb{N}\to\mathbb{N}$ is monotonically increasing.
+  3. (c)
+    Choose a subsequence $(\bm{x}_{m(l)})_{l=1}^{\infty}\subset\mathcal{X}$ of $(\bm{x}_{n})_{n=1}^{\infty}$ satisfying $\lim_{l\to\infty}\mathcal{M}_{\gamma_{m(l)}}^{F_{m(l)},\iota_{C}}(\bm{x}_{m(l)})=0$.
 Then, every cluster point $\bm{x}^{\star}\in C$ of $(\bm{x}_{m(l)})_{l=1}^{\infty}$ is a stationary point of Problem 1.2.
@@
 5.4] at every $\widebar{\bm{x}}\in\mathrm{dom}(\phi)$ (Note: such a $\phi$ includes, e.g., the $\ell_{0}$-pseudonorm).
-In this case, the measure in (11) is extended to $\mathcal{M}_{\gamma}^{F,\phi}(\widebar{\bm{x}})\coloneqq\inf\left\{\left\lVert\frac{\widebar{\bm{x}}-\bm{p}}{\gamma}\right\rVert\mid\bm{p}\in\mathrm{prox}_{\gamma\phi}(\widebar{\bm{x}}-\gamma\bm{v}),\ \bm{v}\in\partial_{F}F(\widebar{\bm{x}})\right\}$.
+In this case, the measure in (11) is extended to $\mathcal{M}_{\gamma}^{F,\phi}(\widebar{\bm{x}})\coloneqq\inf\left\{\left\lVert(\widebar{\bm{x}}-\bm{p})/\gamma\right\rVert\mid\bm{p}\in\mathrm{prox}_{\gamma\phi}(\widebar{\bm{x}}-\gamma\bm{v}),\ \bm{v}\in\partial_{F}F(\widebar{\bm{x}})\right\}$.
 To the problem for finding a stationary point of $F+\phi$, we can apply a modified version of Alg.
-1, in the sense of Thm.
-3.4, by replacing (i) $P_{C}$ in line 5 with $\mathrm{prox}_{\rho^{m}\widetilde{\gamma}}$; and (ii) the condition in line 6 with the condition $(F_{n}+\phi)(\bm{x})\leq(F_{n}+\phi)(\bm{x}_{n})-c\rho^{m}\widetilde{\gamma}\left\lVert\frac{\bm{x}_{n}-\bm{x}}{\rho^{m}\widetilde{\gamma}}\right\rVert$, where $\widetilde{\gamma}\in\mathbb{R}_{++}$ is chosen such that $\phi+\frac{1}{2\widetilde{\gamma}}\left\lVert\cdot\right\rVert^{2}$ is bounded below.
+1 by replacing (i) $P_{C}$ in line 5 with $\mathrm{prox}_{\rho^{m}\widetilde{\gamma}}$; and (ii) the condition in line 6 with the condition $(F_{n}+\phi)(\bm{x})\leq(F_{n}+\phi)(\bm{x}_{n})-c\rho^{m}\widetilde{\gamma}\left\lVert(\bm{x}_{n}-\bm{x})/(\rho^{m}\widetilde{\gamma})\right\rVert^{2}$.
 4 Numerical Experiments on Synthetic Data
@@
 which is $(\theta-1)^{-1}$-weakly convex (e.g., [34]).
-For comparisons, we employed two state-of-the-art models: the model (3) in [1] and the model in [3] with weights $\lambda,\beta\in\mathbb{R}_{++}$:
+For comparisons, we employed two state-of-the-art models: the model (3) in [1] and the following model in [3] with weights $\lambda,\beta\in\mathbb{R}_{++}$:
 $\mathop{\mathrm{minimize}}_{\bm{X}\in\mathbb{R}^{n_{1}\times n_{2}}}\ \left\lVert\bm{y}-\mathcal{A}(\bm{X})\right\rVert_{1}+\lambda(\left\lVert\bm{X}\right\rVert_{\rm nuc}-\beta\left\lVert\bm{X}\right\rVert_{F}),$ (16)
@@
 We generated $\bm{y}\in\mathbb{R}^{m}$ by (1), where each entry of noise $\bm{\epsilon}\in\mathbb{R}^{m}$ was sampled from $\mathcal{N}(0,10^{-6})$ according to [3], the index set $\mathcal{I}_{\rm out}$ of outliers was randomly chosen, and each outlier $\xi_{i}\in\mathbb{R}\ (i\in\mathcal{I}_{\rm out})$ was generated with $\Omega\coloneqq\max_{i=1,2,\ldots,m}{\lvert\mathcal{A}_{i}(\bm{X}^{\star})\rvert}$ from (i) uniform distribution $\mathcal{U}_{[-\Omega,\Omega]}$; and (ii) Cauchy distribution, i.e., $\xi_{i}\coloneqq\Omega\tan(\pi u_{i}/2)$ with $u_{i}\in(-1,1)$ sampled from $\mathcal{U}_{(-1,1)}$.
-Table 1 demonstrates the averaged RMSE and runtime of each model [^5] For (5) with $\ell\coloneqq r_{\theta}^{\rm SCAD}$ in (15), we employed $\theta\in\{2.5,2.7,2.9,3.1\}$.
-For (3), $\lambda\in\{10^{-2},10^{-1},1,2,5\}$, was chosen.
+Table 1 demonstrates the averaged root-mean-square error (RMSE) and runtime of each model [^4] For (5) with $\ell\coloneqq r_{\theta}^{\rm SCAD}$ in (15), we employed $\theta\in\{2.5,2.7,2.9,3.1\}$.
+For (3), $\lambda\in\{10^{-2},10^{-1},1,2,5\}$ was chosen.
 Moreover, stepsize of the subgradient method used in [1] at $n$th iteration was chosen from $0.95^{n}$ or $2\times 0.95^{n}$, where the first choice was suggested by [1] while the second choice achieved better results in some cases in our scenario.
@@
 Moreover, from Table 1, the proposed model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ achieves smaller RMSE than the model (5) with $\ell\coloneqq\lvert\cdot\rvert$ in a shorter runtime; and RMSE of the existing models (3) and (16) are much higher than that of the model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ in particular for $p_{m}\in\{0.3,0.5\}$.
-These observations imply that a weakly convex $r^{\rm SCAD}_{\theta}$ is more robust against outliers, at least in the proposed model (5), than a convex $\lvert\cdot\rvert$ as we expected; and the proposed model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ outperforms the existing models (3) and (16) in terms of recovery accuracy.
+These observations imply that a weakly convex $r^{\rm SCAD}_{\theta}$ is more robust against outliers, at least in the proposed model (5), than a convex $\lvert\cdot\rvert$ as we expected; and the proposed model (5) with $\ell\coloneqq r^{\rm SCAD}_{\theta}$ outperforms the existing models (3) and (16) in terms of recovery accuracy [^5] We note that theoretical recovery guarantees of optimization models involving nonconvex loss functions, e.g., SCAD, have been reported for special cases of RLRMR (2), e.g., robust matrix completion [41].
+However, such a theoretical recovery guarantee of the proposed model (5) is beyond the scope of this paper, and will be addressed in future work.
+.
 Table 1: Averaged RMSE and (averaged runtime [s]) for Problem 1.1
@@
 Our numerical experiments demonstrate that the proposed model with a weakly convex loss function, called SCAD, dramatically improves estimation performance of robust low-rank matrix recovery under severe outliers.
-Finally, we defer further numerical evaluations and complete proofs, including the extensions discussed in Remark 3.5, to an extended manuscript in preparation.
+Finally, we defer further numerical evaluations and complete proofs to an extended manuscript in preparation.